If the numerator of a fraction is increased by 2 and the denominator by 1, it becomes 1. In case the numerator is decreased by 4 and the denominator by 2, it becomes 1/2. Find the fraction.
Answer by mathmadeeasy(2)

Given equations:
1. x+2/y+1 = 1
*multiply both sides by y+1
x+2 = y+1
x = y-1
2. x-4/y-2 = 1/2
2(x-4) = 1(y-2)
2x-8 = y-2
2x = y+6
*insert x = y-1 to equation above
2(y-1)= y+6
2y-2 = y+6
y = 8
So therefore, using x = y-1
x = 8-1
x = 7
Answer:
x = 7
y = 8
Testing:
x = 7 + 2 = 9
y = 8 + 1 = 9
9/9 = 1
x = 7 - 4 = 3
y = 8 - 2 = 6
3/6 = 1/2